# Sparse-IFT: Sparse Iso-FLOP Transformations for Maximizing Training Efficiency

### v3

@@ Abstract
Recent research has focused on weight sparsity in neural network training to reduce FLOPs, aiming for improved efficiency (test accuracy w.r.t training FLOPs). However, sparse weight training often sacrifices accuracy, requiring extended training schedules to attain the accuracy of dense models. In contrast, our approach, Sparse Iso-FLOP Transformations (Sparse-IFT), uses sparsity to improve accuracy while maintaining dense model FLOPs. Using a single hyperparameter (i.e., sparsity level), Sparse-IFTs efficiently replace dense layers, expanding the search space for optimal sparse masks. In addition, dynamic sparse training with Sparse-IFT models effectively navigates this larger sparse mask-weight space, which is evidenced by a spectral analysis using Ramanujan graph properties. Our study reveals a robust correlation among mask topology, weights, and final performance. Notably, without adjusting hyperparameters, replacing dense layers with Sparse-IFT yields significant improvements, such as a +3.5% boost for ResNet-18 on ImageNet and +0.9% for GPT-3 Small on the Open LLM leaderboard. To our knowledge, this is the first work to demonstrate the use of sparsity for improving the accuracy of dense models through a simple-to-use set of sparse transformations. Code is available at: https://github.com/CerebrasResearch/Sparse-IFT.
@@ 1 Introduction
Figure 1: Top-1 Accuracy vs. Training FLOPs for variants of ResNet on ImageNet. Sparse-IFT provides significant accuracy gains across different models and sparsity levels while using the same FLOP budget as its dense counterpart.
@@
Increases in model size and training data have led to many breakthroughs in deep learning (e.g., AlexNet (Krizhevsky et al., 2012), ResNet (He et al., 2016), Transformers (Vaswani et al., 2017), GPT (Radford et al., 2018, 2019), AlphaGo (Silver et al., 2017), etc.). Consequently, computational and memory demands for training and deploying deep neural networks (DNNs) have surged dramatically. To enable the deployment of large models, multiple techniques (e.g., distillation (Hinton et al., 2015), quantization (Han et al., 2015a), pruning (Han et al., 2015b)) have been introduced to reduce inference FLOPs and memory requirements. While these techniques improve inference efficiency (test accuracy w.r.t inference FLOPs), the associated training costs are still prohibitive. Our work focuses on improving the training efficiency (test-accuracy w.r.t training FLOPs) of DNNs through weight sparsity. In recent years, we have witnessed progress in using weight sparsity to reduce training FLOPs (Evci et al., 2020; Liu et al., 2021a; Jayakumar et al., 2020). Frankle & Carbin (2018) show that sparse subnetworks (”lottery tickets”) exist at initialization and can be trained to match dense network accuracy. Dynamic sparse training (DST) methods (Ma et al., 2022; Evci et al., 2020; Liu et al., 2021b; Jayakumar et al., 2020) iteratively adjust sparsity patterns to facilitate the discovery of optimal sparse subnetworks within a single training run. However, they often lag behind dense baselines or require longer training schedules (e.g., 2-5x training steps) to close the gap (Yuan et al., 2021; Tai et al., 2022; Liu et al., 2021a). Our unique contribution focuses on using sparsity to improve a given dense model’s accuracy. We introduce the Sparse Iso-FLOP Transformations (Sparse-IFT), a family of techniques serving as drop-in replacements for dense layers in DNNs.
@@
Sparse-IFTs increase layer representational capacity, facilitating the discovery of optimal sparse subnetworks while maintaining constant FLOPs (i.e., Iso-FLOP). For example, widening a layer with maintained sparsity increases dimensionality without impacting FLOPs; expanding the sparse mask-weight space for more diverse configurations. This enables DST methods to navigate the solution space effectively, potentially finding improved sparse subnetworks for higher accuracy. Drawing inspiration from prior works (Hoang et al., 2023b, a), we analyze the connectivity of Sparse-IFT models as Ramanujan graphs and their impact on performance when trained with DST. All Sparse-IFTs are parameterized by a single hyperparameter, the sparsity level. Figure 1 summarizes ImageNet performance, showing significant accuracy gains with Sparse Wide IFT ResNet variants. Sparse Wide ResNet-18 achieves +3.5% top-1 accuracy at 90% sparsity, surpassing a dense ResNet-34 (74.2%) with 2x fewer FLOPs. These gains result from replacing dense layers with Sparse-IFTs, requiring no changes to training hyperparameters. The main contributions of our work are:
@@
    We showcase the practical value of Sparse-IFT with real-world timings for training on the Cerebras CS-2 (Lie, 2022; Cerebras, 2023) and inference with Neural Magic DeepSparse (NeuralMagic, 2021) using unstructured sparsity. Despite being 2x wider at 75% sparsity with Sparse Wide IFT, we observe minimal compute overhead on both platforms compared to GPUs.
@@ 2 Method
2.1 Training with Dense Matrices is FLOP Inefficient
Modern DNNs are often overparameterized, showing sparsity in features and weights across layers. The Lottery Ticket Hypothesis (Frankle & Carbin, 2018; Chen et al., 2020) suggests sparse DNNs, initialized with an effective sparsity mask (”lottery ticket”), can achieve the same accuracy as dense counterparts. Sparse training methods theoretically enhance efficiency but often yield lower accuracy than dense baselines. This discrepancy may stem from challenges in identifying optimal masks within a single training run. Existing sparse training methods (Jayakumar et al., 2020; Evci et al., 2020; Yuan et al., 2021; Tai et al., 2022; Liu et al., 2021a) invest these FLOP savings into longer training schedules to bridge accuracy gaps, inefficiently requiring more FLOPs than dense baselines for the same target accuracy.
In our work, we take an orthogonal approach and invest these FLOP savings to (1) enhance a layer’s representational capacity and (2) expand its search space, aiming to discover an optimal sparse mask (Ramanujan et al., 2020; Stosic & Stosic, 2021). Larger sparsity-enabled models show potential for improved accuracy, but designing an appropriate architecture is challenging. For instance, achieving performance surpassing ResNet-18 on ImageNet requires careful balance of sparsity and network size. Existing studies explore diverse combinations but often lack FLOP efficiency, requiring multiple iterations for optimal settings and hyperparameter tuning. To address this, we propose the Sparse Iso-FLOP Transformation (Sparse-IFT) family, replacing dense layers with FLOP-equivalent sparse transformations. Notably, Sparse-IFT is parameterized by a single hyperparameter—the sparsity level, simplifying the tuning process.
@@ 2.2 Sparse Iso-FLOP Transformations
For clarity, we explain our method in the context of a fully connected network. Let $\mathcal{N}$ denote a $L$ layered DNN parameterized by $\Theta_{\mathcal{N}}$. Let $\Theta_{\mathcal{N}}\in\{\theta_{1},\ldots,\theta_{L}\}$ denote the parameters of the DNN. The output of the $l$-th layer is defined as: $z_{l}=\sigma(f_{\theta_{l}}(z_{l-1}))$ for some activation function $\sigma$ (e.g., ReLU (Nair & Hinton, 2010)) and feedforward function $f_{\theta_{l}}$. Specifically, let $f_{\theta_{l}}(z_{l-1})=\theta_{l}^{T}z_{l-1}$, where $\theta_{l}\in\mathbb{R}^{D_{in}\times D_{out}}$, $z_{l-1}\in\mathbb{R}^{D_{in}\times B}$ and $B$, $D_{in}$, $D_{out}$ denote the batch-size, input, and output dimensionality of features respectively. The total FLOPs needed for $f_{\theta_{l}}$ are given by $B{\cdot}D_{in}{\cdot}D_{out}$. In Appendix A.1, we detail a straightforward extension to convolutional layers. In the standard setup, the feedforward function $f_{\theta_{l}}$ computes output features through a linear transformation of input features. While theoretically, arbitrary non-linear transformations can be applied, practical implementations often resort to expressing transformations as dense matrix multiplications for efficient GPU support (Nvidia, 2023). We aim to boost DNN training efficiency by enhancing the representational capacity of the feedforward function. Unlike conventional methods that increase capacity by stacking more layers (Lin et al., 2014a), widening (Zagoruyko & Komodakis, 2016), or ensembling (Littwin et al., 2020), our approach introduces unstructured sparsity in weight matrices, achieving the same FLOPs as a dense feedforward function. Let $\Psi_{l}$ denote the set of Sparse Iso-FLOP Transformations (Sparse-IFT) for a particular layer $l$:
@@ 2.3 Members of Sparse-IFT
Increasing the sparsity mask search space with Sparse-IFT is anticipated to enhance training efficiency, as indicated by previous research (Ramanujan et al., 2020; Liu et al., 2022c; Stosic & Stosic, 2021). The likelihood of finding a lottery ticket in a randomly initialized network increases with network width (Ramanujan et al., 2020). Both Liu et al. (2022b) and Stosic & Stosic (2021) show that expanding the search space through increased width or depth improves accuracy. The search space cardinality, defined as the weights a sparse training method can explore, is detailed in Table 1. Sparse Wide, Sparse Parallel, and Sparse Factorized scale with width, parallel branches, and hidden dimension size, respectively. Sparse Doped maintains a constant search space by allocating FLOPs between a low-rank and an unstructured sparse weight matrix. Therefore, dynamic sparse training (DST) becomes crucial for effectively traversing this larger sparse parameter subspace, as discussed in Section 3.
@@ 3 Sparse-IFT and Dynamic Sparse Training
This section explores Sparse-IFT networks trained with DST, showcasing DST’s effectiveness in navigating larger parameter spaces (Huang et al., 2023; Tai et al., 2022). We empirically establish DST’s consistent superiority over static sparse training and then examine the behavior of DST-trained Sparse-IFT models through the lens of Ramanujan graphs. All experiments use the ResNet-18 architecture on CIFAR-100 with published settings (DeVries & Taylor, 2017). Model details and hyperparameters are in Appendix C.1, and results are averaged over 3 seeds.
Importance of DST
@@
Sparse-IFTs employ unstructured sparsity in its transformations. This study investigates the impact of sparse training methods on various Sparse-IFT configurations, focusing on Sparse Wide IFT with sparsity $\in\{50\%,75\%,90\%\}$. We evaluate: random static sparsity, SNIP (Lee et al., 2018), GraSP (Wang et al., 2020a), FORCE(de Jorge et al., 2020), SET (Mocanu et al., 2018), RigL (Evci et al., 2020) and GraNet (Liu et al., 2021a). SET, RigL, and GraNet are DST methods, with SET updating the mask randomly, RigL updating it with gradient information and GraNet incorporating gradual magnitude pruning (Zhu & Gupta, 2017) with RigL. Pruning at Initialization (PaI) methods (e.g., SNIP, GraSP, FORCE) and GraNet increase training FLOPs due to non-uniform sparsity and dense-to-sparse training. We address this by adjusting target sparsity levels to align Sparse-IFT training FLOPs with the dense baseline (see Appendix A.2). Results in Table 2 indicate that DST methods consistently outperform static sparsity, with improvements persisting at higher sparsity levels. However, in Iso-FLOP scenarios, PaI methods underperform, likely due to disruptive sparsity imbalances affecting information propagation. Sparse-IFTs expand the sparse mask-weight space $\propto$ sparsity, benefiting DST methods in thorough exploration and exploitation within this space. RigL is chosen as the sparse training method for simplicity in all experiments, despite GraNet performing the best.
3.1 Spectral Analysis of Sparse-IFT trained with DST

### v4

@@ Abstract
Recent research has focused on weight sparsity in deep neural network training to reduce FLOPs, aiming for improved efficiency (test accuracy w.r.t training FLOPs). However, sparse weight training often compromises accuracy, requiring extended training schedules to attain the accuracy of dense models. In contrast, our approach, Sparse Iso-FLOP Transformations (Sparse-IFT), uses sparsity to improve accuracy while maintaining dense model FLOPs. Using a single hyperparameter (i.e., the sparsity level), Sparse-IFTs efficiently replace dense layers, expanding the search space for optimal sparse masks. In addition, dynamic sparse training (DST) with Sparse-IFT models effectively navigate this larger sparse mask-weight space, which is evidenced by a spectral analysis using Ramanujan graph properties. Our study reveals a robust correlation among mask topology, weights, and final performance. Notably, without adjusting any training hyperparameters, replacing dense layers with Sparse-IFT yields significant improvements, such as a +3.5% boost for ResNet-18 on ImageNet and +0.9% for GPT-3 Small on the Open LLM leaderboard. To the best of our knowledge, this is the first work to demonstrate the use of sparsity for improving the accuracy of dense models through a set of simple-to-use sparse transformations. Code is available at: https://github.com/CerebrasResearch/Sparse-IFT.
@@ 1 Introduction
[…]
@@
Increases in model size and training data have led to many breakthroughs in deep learning (e.g., AlexNet (Krizhevsky et al., 2012), ResNet (He et al., 2016), Transformers (Vaswani et al., 2017), GPT (Radford et al., 2018, 2019), AlphaGo (Silver et al., 2017), etc.). Consequently, computational and memory demands for training and deploying deep neural networks (DNNs) have surged dramatically. To enable the deployment of large models, multiple techniques (e.g., distillation (Hinton et al., 2015), quantization (Han et al., 2015a), pruning (Han et al., 2015b)) have been introduced to reduce inference FLOPs and memory requirements. While these techniques improve inference efficiency (test accuracy w.r.t inference FLOPs), the associated training costs are still prohibitive. Our work focuses on improving the training efficiency (test-accuracy w.r.t training FLOPs) of DNNs through weight sparsity. In recent years, we have witnessed progress in using weight sparsity to reduce training FLOPs (Evci et al., 2020; Liu et al., 2021a; Jayakumar et al., 2020). Frankle & Carbin (2018) show that sparse subnetworks (“lottery tickets”) exist at initialization and can be trained to match dense network accuracy. Dynamic sparse training (DST) methods (Ma et al., 2022; Evci et al., 2020; Liu et al., 2021b; Jayakumar et al., 2020) iteratively adjust sparsity patterns to facilitate the discovery of optimal sparse subnetworks within a single training run. However, they often lag behind dense baselines or require longer training schedules (e.g., 2-5x training steps) to close the gap (Yuan et al., 2021; Tai et al., 2022; Liu et al., 2021a). Our unique contribution focuses on using sparsity to improve a given dense model’s accuracy. We introduce the Sparse Iso-FLOP Transformations (Sparse-IFT), a family of techniques serving as drop-in replacements for dense layers in DNNs.
@@
Sparse-IFTs increase layer representational capacity, facilitating the discovery of optimal sparse subnetworks while maintaining constant FLOPs (i.e., Iso-FLOP). For example, widening a layer with maintained sparsity increases dimensionality without impacting FLOPs; expanding the sparse mask-weight space for more diverse configurations. This enables DST methods to navigate the search space effectively, potentially finding improved sparse subnetworks for higher accuracy. Drawing inspiration from prior works (Hoang et al., 2023b, a), we analyze the connectivity of Sparse-IFT models as Ramanujan graphs and their impact on performance when trained with DST. All Sparse-IFTs are parameterized by a single hyperparameter, the sparsity level. Figure 1 summarizes ImageNet performance, showing significant accuracy gains with Sparse Wide IFT ResNet variants. Sparse Wide ResNet-18 achieves +3.5% top-1 accuracy at 90% sparsity, surpassing a dense ResNet-34 (74.2%) with 2x fewer FLOPs. These gains result from replacing dense layers with Sparse-IFTs, requiring no changes to training hyperparameters. The main contributions of our work are:
@@
    We showcase the practical value of Sparse-IFT with real-world timings for training on the Cerebras CS-2 (Lie, 2023) and inference with Neural Magic DeepSparse (NeuralMagic, 2021) using unstructured sparsity. Despite being 2x wider at 75% sparsity with Sparse Wide IFT, we observe minimal compute overhead on both platforms compared to GPUs.
@@ 2 Method
Training with Dense Matrices is FLOP Inefficient
Modern DNNs are often overparameterized, showing sparsity in features and weights across layers. The Lottery Ticket Hypothesis (Frankle & Carbin, 2018; Chen et al., 2020) suggests sparse DNNs, initialized with an effective sparsity mask (“lottery ticket”), can achieve the same accuracy as dense counterparts. Sparse training methods theoretically enhance efficiency but often yield lower accuracy than dense baselines. This discrepancy may stem from challenges in identifying optimal masks within a single training run. Existing sparse training methods (Jayakumar et al., 2020; Evci et al., 2020; Yuan et al., 2021; Tai et al., 2022; Liu et al., 2021a) invest these FLOP savings into longer training schedules to bridge accuracy gaps, inefficiently requiring more FLOPs than dense baselines for the same target accuracy.
In our work, we take an orthogonal approach and invest these FLOP savings to (1) enhance a layer’s representational capacity and (2) expand its search space, aiming to discover an optimal sparse mask (Ramanujan et al., 2020; Stosic & Stosic, 2021). Larger sparse models show potential for improved accuracy, but designing an appropriate architecture is challenging. For instance, achieving performance surpassing ResNet-18 on ImageNet requires careful balance of sparsity and network size. Existing studies explore diverse combinations but often lack FLOP efficiency, requiring multiple iterations for optimal settings and hyperparameter tuning. To address this, we propose the Sparse Iso-FLOP Transformation (Sparse-IFT) family, replacing dense layers with FLOP-equivalent sparse transformations. Notably, Sparse-IFT is parameterized by a single hyperparameter—the sparsity level, simplifying the tuning process.
@@ 2.1 Sparse Iso-FLOP Transformations
For clarity, we explain our method in the context of a fully connected network. Let $\mathcal{N}$ denote a $L$ layered DNN parameterized by $\Theta_{\mathcal{N}}$. Let $\Theta_{\mathcal{N}}\in\{\theta_{1},\ldots,\theta_{L}\}$ denote the parameters of the DNN. The output of the $l$-th layer is defined as: $z_{l}=\sigma(f_{\theta_{l}}(z_{l-1}))$ for some activation function $\sigma$ (e.g., ReLU (Nair & Hinton, 2010)) and feedforward function $f_{\theta_{l}}$. Specifically, let $f_{\theta_{l}}(z_{l-1})=\theta_{l}^{T}z_{l-1}$, where $\theta_{l}\in\mathbb{R}^{D_{in}\times D_{out}}$, $z_{l-1}\in\mathbb{R}^{D_{in}\times B}$ and $B$, $D_{in}$, $D_{out}$ denote the batch-size, input, and output dimensionality of features respectively. The total FLOPs needed for $f_{\theta_{l}}$ are given by $B{\cdot}D_{in}{\cdot}D_{out}$. In Appendix A.1, we detail a straightforward extension to convolutional layers. In the standard setup, the feedforward function $f_{\theta_{l}}$ computes output features through a linear transformation of input features. While theoretically, arbitrary non-linear transformations can be applied, practical implementations often resort to expressing transformations as dense matrix multiplications for efficient GPU support (Nvidia, 2023). We aim to boost DNN training efficiency by enhancing the representational capacity of the feedforward function. Unlike conventional methods that increase capacity by stacking more layers (Lin et al., 2014a), widening (Zagoruyko & Komodakis, 2016), or ensembling (Littwin et al., 2020), our approach introduces unstructured sparsity in weight matrices, achieving the same FLOPs as a dense feedforward function.
Let $\Psi_{l}$ denote the set of Sparse Iso-FLOP Transformations (Sparse-IFT) for a particular layer $l$:
@@ 2.2 Members of Sparse-IFT
Increasing the sparsity mask search space with Sparse-IFT is anticipated to enhance training efficiency, as indicated by prior works (Ramanujan et al., 2020; Liu et al., 2022c; Stosic & Stosic, 2021). The likelihood of finding a lottery ticket in a randomly initialized network increases with network width (Ramanujan et al., 2020). Both Liu et al. (2022b) and Stosic & Stosic (2021) show that expanding the search space through increased width or depth improves accuracy. The search space cardinality, defined as the weights a sparse training method can explore, is detailed in Table 1. Sparse Wide, Sparse Parallel, and Sparse Factorized scale with width, parallel branches, and hidden dimension size, respectively. Sparse Doped maintains a constant search space by allocating FLOPs between a low-rank and an unstructured sparse weight matrix. Therefore, DST becomes crucial for effectively traversing this larger parameter subspace, as discussed in Section 3.1.
3 Sparse-IFT Ablation Studies
In this section, we present a comprehensive analysis of Sparse-IFT networks, focusing on their training methodologies and design considerations. First, we compare static sparse training with DST, highlighting DST’s superior performance in handling larger parameter spaces through empirical results using the ResNet-18 architecture on CIFAR-100. Then, we explore critical design aspects of Sparse-IFT, including the role of non-linearities, and the benefits of dynamic unstructured sparsity over structured sparsity. Finally, we evaluate the efficacy of DST by comparing it against densely trained Sparse-IFT models.
3.1 Impact of Sparse Training Techniques
@@
This section provides a comparative analysis of Sparse-IFT networks trained with two classes of methods: static sparse training and DST. The focus is on demonstrating DST’s effectiveness in navigating larger parameter spaces, as evidenced by previous research (Huang et al., 2023; Tai et al., 2022). Our empirical results consistently show DST’s superiority over static sparse training. All experiments utilize the ResNet-18 architecture on CIFAR-100 with published settings (DeVries & Taylor, 2017). Detailed model information and hyperparameters are available in Appendix C.1, and all results are averaged over 3 seeds.
@@
Sparse-IFTs employ unstructured sparsity in its transformations. This study investigates the impact of sparse training methods on various Sparse-IFT configurations, focusing on Sparse Wide IFT with sparsity $\in\{50\%,75\%,90\%\}$. In Table 2, we evaluate: random static sparsity, SNIP (Lee et al., 2018), GraSP (Wang et al., 2020a), FORCE (de Jorge et al., 2020), SET (Mocanu et al., 2018), RigL (Evci et al., 2020) and GraNet (Liu et al., 2021a). SET, RigL, and GraNet are DST methods, with SET updating the mask randomly, RigL updating it with gradient information and GraNet incorporating gradual magnitude pruning (Zhu & Gupta, 2017) with RigL. Pruning at Initialization (PaI) methods (e.g., SNIP, GraSP, FORCE) and GraNet increase training FLOPs due to non-uniform sparsity and dense-to-sparse training. We address this by adjusting target sparsity levels to align Sparse-IFT training FLOPs with the dense baseline (see Appendix A.2). In Iso-FLOP scenarios, PaI methods underperform because they heavily prune parameter-rich layers to match target sparsity levels, leading to layer-collapse and poor gradient flow. Furthermore, DST methods consistently outperform static sparsity, with improvements persisting at higher sparsity levels. Sparse-IFTs expand the sparse mask-weight space $\propto$ sparsity, benefiting DST in thorough exploration and exploitation within this space. While RigL and GraNet attain similar performance, RigL is chosen as the sparse training method for simplicity in all experiments.
3.2 Assessing the Effects of Architecture Variations
This section analyzes different design considerations for Sparse-IFT by first, exploring the role of non-linearities in enhancing representational capacity. Then, the advantage of training with dynamic unstructured sparsity over structured is investigated. Next, we compare between densely and sparsely trained Sparse-IFT models. Finally, by applying top-performing Sparse-IFTs to efficient vision models, these insights contribute to a synthesized framework.
Importance of Using Non-Linear Activations
For some of the Sparse-IFT members, we draw inspiration from linear overparameterization methods, which fold the feedforward function into a dense matrix post-training (Ding et al., 2021b, a; Guo et al., 2020a; Ding et al., 2019). Our method enhances representational capacity through an Iso-FLOP transformation without increasing training FLOPs. Maintaining original dense FLOP levels eliminates the need for post-training modifications, enabling efficient inference and incorporation of non-linearities (i.e., ReLU) in Sparse-IFT. Experiments on ResNet-18 on CIFAR-100 show notable accuracy gains across all sparsity levels with non-linear activations. For example, at 90% sparsity, using non-linearities in Sparse Factorized IFT yields a 1.8% accuracy increase over the dense baseline, in contrast to a 0.5% decrease without non-linearities. These findings extend to all Sparse-IFT members (see Appendix C.2 for details). The accuracy improvements at all sparsity levels highlight the effectiveness of incorporating non-linear activations in Sparse-IFT.
Unstructured vs. Structured Sparsity
We compare dynamic unstructured and structured sparsity using Sparse-IFT. Unstructured sparsity explores all mask variations, but most hardware accelerators do not support unstructured sparse acceleration. Prior works have investigated structured sparsity, such as low-rank and block-sparse matrices, for wall-clock speed-ups (Khodak et al., 2020; Chen et al., 2021b; Hubara et al., 2021; Dao et al., 2022). We explore structured sparsity through Iso-FLOP configurations with Sparse Wide IFT, employing low-rank factorization and N:M sparsity for GPU acceleration. In Figure 3 (left plot), we compare dynamic unstructured sparsity with N:M transposable structured sparsity (Hubara et al., 2021) using Sparse-IFT. The latter demonstrates improvements over the dense baseline at 75% and 90% sparsity levels. Results also indicate that N:M block sparsity outperforms low-rank factorization (see Appendix C.3.3). However, unstructured sparsity still gives the highest gains, as N:M sparsity has reduced mask diversity in block-sparse matrices (Hubara et al., 2021), therefore, we adopt unstructured sparsity in all subsequent experiments.
Sparse-IFT ResNet-18
We assess all Sparse-IFT family members with ResNet-18 on CIFAR-100 across different sparsity levels. The middle plot of Figure 3, highlights the best accuracy achieved by each Sparse-IFT member. All members exhibit substantial accuracy improvements compared to the dense baseline (77%), using the same FLOPs. Sparse Wide consistently performs the best, while Sparse Doped is the only member not gaining accuracy at higher sparsity. This is attributed to Sparse Doped maintaining constant search space by distributing FLOPs between low-rank and unstructured sparse matrices (see Table 1), leading to a decrease in active weights in the unstructured matrix. In Appendix C.3.1, we compare Sparse-IFT against other DST baselines under the same training efficiency setup by extending the training steps, showing Sparse-IFT outperforms them significantly at $s\in\{50\%,75\%,90\%\}$. Since, Sparse Parallel and Sparse Wide perform the best across ablations, we use these two IFTs for the main experiments.
Sparse-IFT vs. Dense Overparametrization